OpenStudy (anonymous):

How do I find the systems of these and graph them? 3x + 6y - 12 = 0 x + 2y = 8

OpenStudy (anonymous):

these are two straight lines plot values of x say -4 to 4 and their y values where they intersect gives you the coordivates of the solution of the system

OpenStudy (anonymous):

Can you go by steps?

hero (hero):

3x + 6y - 12 = 0 x + 2y = 8 Solve both equations for y: \[y = \frac{12-3x}{6}\]\[y = \frac{8-x}2 \] Set y = y and solve for x: \[\frac{12-3x}{6} = \frac{8-x}{2}\] \[2(12-3x) = 6(8-x)\] \[24 - 6x = 48 - 6x\] No Solution

OpenStudy (anonymous):

do u want to solve them firsr ?

OpenStudy (anonymous):

Thanks the steps really helped a lot. :)